OpenStudy (anonymous):

STUDY GUIDE HELP Use the Law of Sines to find the missing side of the triangle.Find the measure of b, given m∠A = 38°, m∠B = 74°, and a = 31.

OpenStudy (precal):

use sinA/a=sinB/b to find b

OpenStudy (precal):

180-38-74 gives you angle C recall all three angles must add up to 180 degrees

OpenStudy (precal):

use sinA/a=sinC/c to find c

OpenStudy (anonymous):

@precal so would it be sin(38)/a?

OpenStudy (anonymous):

=sin(74)/a?

OpenStudy (precal):